OpenStudy (joehoffy):

Which of the following correctly simplifies the expression 2^3*3^0/5?

OpenStudy (joehoffy):

\[{2^3*3^0\over5}^2\]

OpenStudy (joehoffy):

thats the expression

OpenStudy (joehoffy):

and it is a multiple choice question

OpenStudy (anonymous):

Well, 3^0 is equal to 1 so you have \[(\frac{ 2^3 \times1}{ 5 })^2\]

OpenStudy (anonymous):

You understand, right?

OpenStudy (joehoffy):

yes

OpenStudy (anonymous):

8/5

OpenStudy (anonymous):

@John4m2 Do not give direct answers!

OpenStudy (joehoffy):

john its a multiple choice question and thats not an answer :)

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

It''s ok.

OpenStudy (anonymous):

So, if you have \[(\frac{ 2^3 \times1^2}{ 5 })\] your answer is...

OpenStudy (joehoffy):

waith tho there are 2 answers with that, did you look at the picture?

OpenStudy (anonymous):

2^3 x 3^0/5 but 3^0=1 2^3 x 1/5 (8 x 1/5)^1

OpenStudy (anonymous):

When you square it, you get \[\frac{ 2^6 \times1^2}{ 5^2 }\]

OpenStudy (joehoffy):

so that would be A!

OpenStudy (anonymous):

Yes!

OpenStudy (joehoffy):

thank you, can you help me with some more?

OpenStudy (joehoffy):

@Lovelarap

OpenStudy (anonymous):

Sure! Sorry, was do my homework...

OpenStudy (joehoffy):

haha its ok.

OpenStudy (anonymous):

ok, next question.

OpenStudy (joehoffy):

ok ill tag u.

OpenStudy (anonymous):

OK